HUBBLE SPACE TELESCOPE DAILY REPORT #935 PERIOD COVERED: 0800 EST Daily Status Report as of 204/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 4893 (Orbital Motions of Close Binary Stars Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system GLIESE-568 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. 1.2 Completed Three Sets of PC/WFC 4809 (Cycle 2 : WF/PC PSF Calibration Proposal) The Planetary Camera was used to make three point spread function calibration images of standard star HD144470. The OSS monitored the observations and assessed the data quality. All three observations were displayed. They all looked okay, but the first observation had a cosmic ray hit across the image's point spread function. 1.3 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make three Earth Calibration Flat images to monitor a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The first two observations were okay, and the third was reported to be saturated. 1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.5 Completed PC 4521 (Asteroid Shape and Duplicity - Cycle 3 High) The Planetary Camera was used to make three snapshot images of the large asteroid AST-709-FRIN to look for possible satellites, shapes, and albedo variations. All three observations were received and displayed, but they showed no target. It would appear that the telescope was not correctly pointed for this moving target. 1.6 Completed FOS/RD/BL 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make two spectrographic observations of quasar 0052+2509 and 0145+3839 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The observations was received in real time. The science observations were reported to show a good spectrum. 1.7 Completed FOS/BL 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle 3 High) The Faint Object Spectrograph was used to make two observations of quasar 1701+6102 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The observations were executed as scheduled, and the spectral observations were reported to look okay. 1.8 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were not monitored closely, but were seen executing (i.e., shutter open) as expected. 1.9 Completed FOS/BL/RD 4976 (UV Rotation Light Curves for Pluto, and Charon's UV Spectrum - Split) The Faint object Spectrograph was used to make a TALED exposure subsequent to other spectral observations of the planet Pluto and its moon Charon. After a corrective SIP maneuver was uplinked to the spacecraft, spectral observations were made of the planet Pluto. The observations were executed in fine lock. No problems were reported. 1.10 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. daises and measles), and the possible return of quantum efficiency hysteresis. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 2 Successful : 2 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of PC 4809 (Cycle 2: WF/PC PSF Calibration Proposal) 3.2 Complete FGS 2565 (Linking the HIPPARCOS to the Extragalactic Reference Frame Part 1 of 6) 3.3 Complete Two Sets of FOS/RD 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle 3 High) 3.4 Complete Four Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Deltas Flats) 3.5 Complete GHRS/2 5012 (Physical and Chemical Processes in Dense Interstellar Clouds) 3.6 Complete Two Sets of GHRS/2 3941 (Physical Parameters in the Local Interstellar Medium - Cycle 2) 3.7 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Narrow. 2) 3.8 Complete Eight Sets of PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) 3.9 Complete FOC/96 3870 (Imaging of M-31 Group Galaxies - Cycle 2) 3.10 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.11 Complete FGS 5068 (Linking the HIPPARCOS to the Extragalactic Reference Frame Part 6 of 6) 3.12 Complete FOS /BL 4328 (The Companions of Dwarf Carbon Stars: Cycle 3 Medium) 3.13 Complete Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.14 Complete GHRS/2 3935 (Sulfur Near IO - Cycle 2) 3.15 Complete FOC/96 4800 (The Stellar Content of Wolf- Rayet Galaxies (FOC Images)) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-